The measurements of a temperature sensor (in °F) are provided. Is the set of measurements well described by a normal distribution? Calculate the R^2 and R^2 adjusted values. Show all the calculations.

Order Temperature (F)
1 118
2 119
3 120
4 121
5 121
6 122
7 123
8 123
9 124
10 124
11 124
12 124
13 125
14 125
15 125
16 125
17 125
18 126
19 127
20 128
21 128
22 129
23 129
24 129
25 130
26 131
27 131
28 131
29 133
30 133
31 136
32 137
33 137
34 140
35 140
36 141
37 142
38 142
39 151
40 160
Answer #1

I have used the R software to plot the values in order to check if its normal. I will attach the plot and the commands along with its output below.

>x=c(118,119,120,121,121,122,123,123,124,124,124,124,125,125,125,125,125,126,127,128,128,129,129,129,130,131,131,131,133,133,136,137,137,140,140,141,142,142,151,160)
> qqnorm(x)
> qqline(x)

120 130 Sample Quantiles 140 150 160 Y T Theoretical Quantiles Normal Q-Q Plot
If most of the values lie on the line, the set of values are said to follow a normal distribution. Since most of our values does not pass through our line, we can consider it not to be normal.

To find the R2 and R2 adjusted values,

Let us consider X as the order and Y as Fahrenheit and find the regression equation.

The regression equation is Ý = b +bx

Where, S.

(2*19) - h = 09

here, 40 Χ;)2 Sty = Σχ? (Σ 40

S) - ΣΧή - Σ, Χ. - Σκαι 40

2 and y = 2 40 40

After substituting the values, we get Sxx=5330 and Sxy=3719.5

T = 20.5 and y = 129.975

b1 = 3719.5 -= 0.6978 5330

bo = 129.975 - (0.6978 * 20.5) = 115.6701

The regression equation becomes \hat{Y}=115.6701+0.6978X

R² = 1- SSE Syy;

where SSE=sum of squares of errors = Σ(vi – 9:02 =503.3502

S_{yy}=\sum_{i=1}^{40}Y_{i}^{2}-\frac{(\sum_{i=1}^{40}Y_{i})^{2}}{40}=3098.975

R^{2}=1-\frac{503.3502}{3098.975}=0.8376

n-1 Rad = 1 n (1 - R2

Here, k-number of independent variables = 1

R^{2}_{adj}=1-\frac{39}{38}(1-0.8376)=0.8333

Therefore, our model explains 83.33% of variability.